a) How many micrometres are represented between a 10th division and the next? Show how you have arrived at your answer. (2 marks)
1
Expert's answer
2021-07-21T10:31:02-0400

complete question

solution



"1\space division =\frac{1}{100}=0.01mm"

between 10 and the next number of division =10

"1 \space division=0.01mm\\
10\space division=0.01\times 10=0.1mm"

"=100\u03bcm=(1mm=10^3\u03bcm)"